# Air-HOLP: Adaptive Regularized Feature Screening for High Dimensional Data

### v1

@@ Abstract
Handling high-dimensional datasets presents substantial computational challenges, particularly when the number of features far exceeds the number of observations and when features are highly correlated. A modern approach to mitigate these issues is feature screening. In this work, the High-dimensional Ordinary Least-squares Projection (HOLP) feature screening method is advanced by employing adaptive ridge regularization. The impact of the ridge penalty on the Ridge-HOLP method is examined and Air-HOLP is proposed, a data-adaptive advance to Ridge-HOLP where the ridge-regularization parameter is selected iteratively and optimally for better feature screening performance. The proposed method addresses the challenges of penalty selection in high dimensions by offering a computationally efficient and stable alternative to traditional methods like bootstrapping and cross-validation. Air-HOLP is evaluated using simulated data and a prostate cancer genetic dataset. The empirical results demonstrate that Air-HOLP has improved performance over a large range of simulation settings. We provide R codes implementing the Air-HOLP feature screening method and integrating it into existing feature screening methods that utilize the HOLP formula.
Keywords: Dimensionality reduction, Regularization, Ridge regression, Sure screening
@@ 1 Introduction
Feature screening is a dimensionality reduction technique that aims to efficiently eliminate numerous non-true features while retaining all true features. Feature screening is often used prior to model selection to handle the curse of dimensionality (Liu and Li, 2020; Liu et al., 2015; Fan et al., 2009). In this paper, we build on the existing High-dimensional Ordinary Least-squares Projection (HOLP) method as introduced in Wang and Leng (2016), by incorporating adaptive ridge regularization. We analyze the impact of the ridge penalty on Ridge-HOLP (Wang and Leng, 2016) and propose Air-HOLP, a method that utilizes a data-adaptive approach for selecting the ridge penalty, which shows remarkable gains in feature screening performance. Additionally, we provide R codes implementing the Air-HOLP feature screening method and integrating it into the group HOLP (Qiu and Ahn, 2020) method which utilizes the HOLP formula. The codes are made available on GitHub at https://github.com/Logic314/Air-HOLP.git.
@@ 2.2 HOLP and Ridge-HOLP
where $r$ is the ridge penalty parameter and $\mathds{I}_{n}$ denotes the identity matrix.
The Ridge-HOLP formula in Equation (3) is equal to the standard ridge formula but is more efficient when $p>n$ as the computational complexity of Ridge-HOLP is $O(n^{3}+n^{2}p)$, while the standard ridge formula is $O(p^{3}+p^{2}n)$. This is because Ridge-HOLP requires an ($n\times p$) by ($p\times n$) multiplication $O(n^{2}p)$ and an ($n\times n$) matrix inverse $O(n^{3})$, while standard ridge requires a ($p\times n$) by ($n\times p$) multiplication $O(p^{2}n)$ and a ($p\times p$) matrix inverse $O(p^{3})$. When $n$ is close to $p$, Wang and Leng (2016) recommend using Ridge-HOLP with a fixed penalty parameter $r=10$.
The penalty in the Ridge-HOLP formula in Equation (3) also helps to handle multicollinearity. When features are highly correlated, then estimated regression coefficients have large variances (Hoerl and Kennard, 1970), which leads to poor ranking accuracy and thus lower screening performance. The penalty in Ridge-HOLP reduces the estimation sensitivity. However, selecting an appropriate penalty parameter $r$ is crucial to balance the bias-variance trade-off and to maximize the screening performance of Ridge-HOLP.
@@
2.3 Challenges of Penalty Selection in High Dimensions
Several methods exist to choose the ridge penalty. Common approaches include bootstrapping (Delaney and Chatterjee, 1986), cross-validation (Allen, 1971, 1974) and generalized cross-validation (Golub et al., 1979). However, such sub-sampling in the Ridge-HOLP method is computationally expensive due to the repeated computation of the Ridge-HOLP formula. To address this, a more efficient formula called multiridge was proposed by van de Wiel et al. (2021) for applying fast cross-validation in ridge regression. The multiridge formula avoids redundant computations of the ridge formula when applied to multiple values of the penalty parameter $r$, reducing the heavy computations to be repeated only $k$ times in k-fold cross-validation. However, k-fold cross-validation is non-deterministic and can therefore introduce instability in feature screening as different data splits may lead to different screened features. Although repeated cross-validation can mitigate this instability, it comes at the cost of increased computational burden (Martinez et al., 2011).
To avoid the computational burden and instability associated with sub-sampling, an alternative approach is to use a closed-form formula for selecting the penalty parameter $r$. Examples of this alternative approach include Hura Ahmad et al. (2006); Alkhamisi and Shukur (2007); Batah et al. (2008); Dorugade and Kashid (2010); Hoerl and Kennard (1970); Hoerl et al. (1975); Lawless (1976); Kibria (2003); Nomura (1988). However, the formulas proposed in these papers require $p$ to be smaller than $n$. Cule and De Iorio (2013) addressed this limitation by extending the Hoerl et al. (1975) approach to handle $p>n$. Nevertheless, their method involves computing the matrix $\boldsymbol{X}^{\top}\boldsymbol{X}$ and its eigendecomposition, resulting in a time complexity of $O(p^{3}+p^{2}n)$.
@@ 3 The Air-HOLP method
Air-HOLP is an adaptive iterative variant of Ridge-HOLP which efficiently selects $r$ while also satisfying the sure screening property. Theorem $3$ of Wang and Leng (2016) and Wang et al. (2021) mentions that for the Ridge-HOLP to achieve sure screening, the penalty parameter $r$ must satisfy $r=o(n^{1-(5/2)\tau-\kappa})$, where $01-7.5\tau-2\kappa-\nu>$ 0, and $\tau,\kappa,\nu>0$. It is worth noting that any $r=cn^{a}$ where $0\leq a\leq 0.5$ and $c\geq 0$ satisfies the required condition on $r$. Thus, to ensure that we achieve the sure screening property, we confine the penalty parameter to the range $[0,c\sqrt{n}]$ for some positive constant $c$ and choose $c=1000$ if not otherwise mentioned. To balance the trade-off between bias and variance we solve
@@
To estimate the unknown expected response $\boldsymbol{y}_{0}$ in (5) we use Algorithm 1, which consists of a feature screening step and then uses the screened features to obtain the estimated expected response $\boldsymbol{\tilde{y}}_{r_{0}}$.
Algorithm 1 Estimation of the expected response vector
$\boldsymbol{X}$, $\boldsymbol{y}$, $r_{0}$, $m$
$\boldsymbol{\tilde{y}}_{r_{0}}$
Step 1: For a given initial penalty parameter $r_{0}$, compute the initial Ridge-HOLP estimator $\boldsymbol{\hat{\beta}}_{r_{0}}=\boldsymbol{X}^{\top}(\boldsymbol{X}\boldsymbol{X}^{\top}+r_{0}\mathds{I}_{n})^{-1}\boldsymbol{y}$.
Step 2: Rank the features in $\boldsymbol{X}$ by the absolute values of their estimated coefficients $|\boldsymbol{\hat{\beta}}_{r_{0}}|$. Then screen the top $m$ features.
Step 3: Fit a regression model utilizing the screened features from Step 2. Then, output $\boldsymbol{\tilde{y}}_{r_{0}}$, the fitted response by the model.
Notably, the third step in Algorithm 1 does not detail a specific method for fitting the model using the screened features from the second step. For simplicity, we employ ordinary least squares regression in our empirical research below. A given penalty parameter $r_{i}$ is updated through
@@
Air-HOLP iteratively updates the penalty parameter $r$ until $|r_{i}-r_{i-1}|<\delta r_{i}$ for some small $\delta>0$, i.e. until the absolute relative error is less than $\delta$ or until a predetermined maximum number of iterations is reached.
@@
$\begin{split}r_{i+1}=\operatorname*{arg\,min}_{0\leq r\leq c\sqrt{n}}{\boldsymbol{y}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}^{2}(\boldsymbol{D}_{n}+r\mathds{I}_{n})^{-2}\boldsymbol{U}^{\top}\boldsymbol{y}-2\boldsymbol{\tilde{y}}_{r_{i}}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}(\boldsymbol{D}_{n}+r\mathds{I}_{n})^{-1}\boldsymbol{U}^{\top}\boldsymbol{y}},\end{split}$ (8)
which is solved by finding the root of its derivative by Newton’s method.
@@ 4 Simulation Study
In this section, we empirically demonstrate the good performance of Air-HOLP. We primarily compare Air-HOLP’s screening performance to Ridge-HOLP with fixed penalty parameter $r=10$ as recommended by Wang and Leng (2016). While we initially also include Sure Independence Screening (SIS) in our comparison, our main focus is to showcase how Air-HOLP advances Ridge-HOLP by using an adaptive penalty parameter in the more general case when features are correlated. In all simulations, Air-HOLP is implemented with an initial penalty parameter $r_{0}=10$, selected penalty $\hat{r}\in[0,1000\sqrt{n}]$, and screening threshold $m=\lceil n/\log(n)\rceil$. The full simulation results and the code implementing Air-HOLP are both made available on GitHub at https://github.com/Logic314/Air-HOLP.git.
@@ 4.1 Simulation Setup
We generate $3200$ unique $\boldsymbol{X}$ samples across $64$ distinct simulation settings, with $50$ samples per setting. For each distinct $\boldsymbol{X}$, we generate $250$ $\boldsymbol{y}$’s across $25$ distinct simulation settings. This generates $800,000$ datasets [$\boldsymbol{y},\boldsymbol{X}$], encompassing $1,600$ distinct simulation settings, with $500$ samples per setting. The specific simulation settings employed are as follows.
We generate the design matrix $\boldsymbol{X}\in\operatorname{mat}(n,p)$ from a multivariate normal distribution with covariance matrix $\boldsymbol{\Sigma}=(1-\rho)\mathds{I}_{p}+\rho$ for varying parameters $n$, $p$, and $\rho$.
@@ 4.3 Results
Here we analyze the results from all $1600$ simulation settings to assess the overall performance and comparison of the competing methods. Our analysis begins by demonstrating the effect of correlation on Air-HOLP, Ridge-HOLP, and SIS. Then, we focus on comparing Air-HOLP to Ridge-HOLP aiming to provide a fair and balanced narrative of the overall performance and comparison between them by showing results for representative settings that highlight differences but we conciously refrained from showing settings that give maximal difference between the methods.
Figure 1: Line plots comparing the performances of Air-HOLP, Ridge-HOLP and SIS for different levels of correlation in two different simulation settings. For Ridge-HOLP, four values for the penalty parameter are compared: $r=10,50,200,600$.
@@
In addition, Figure 1 also demonstrates the advantage of adaptive regularization by showcasing two settings were the optimal penalty parameter differ significantly. The lower penalty parameters ($r=10$ and $r=50$) performs well in Figure 1(a) but not as well in Figure 1(b), whereas the higher penalty parameters ($r=200$ and $r=600$) show the opposite trend. Air-HOLP, on the other hand, consistently performs well in both settings.
Figure 2: Sure Screening Probability Heatmaps. The left panel shows Air-HOLP, the middle panel shows Ridge-HOLP (with $r=10$), and the right panel shows the difference between Air-HOLP and Ridge-HOLP (Sure Screening Probability of Air-HOLP minus Ridge-HOLP).
We now shift the focus to comparing Air-HOLP and Ridge-HOLP with the recommended penalty parameter of $r=10$, as suggested by Wang and Leng (2016). Across all $1600$ simulation settings, the difference in sure screening probability between Air-HOLP and Ridge-HOLP ranged from $-0.02$ to $0.60$. In $47\%$ of the settings, the sure screening probabilities for Air-HOLP and Ridge-HOLP were equal, typically occurring when both probabilities were either 00 or $1$. In $45\%$ of the settings, Air-HOLP’s sure screening probability was higher than Ridge-HOLP’s, while in $8\%$ of the settings, it was lower. To better understand the factors influencing this gap in performance, we analyze the impact of $n,p,p_{0}$, and $R^{2}$ on feature screening performance of Air-HOLP and Ridge-HOLP.
@@
(a) The logarithm of the Sure Screening Threshold vs. Theoretical $R^{2}$.
(b) The logarithm of the Sure Screening Threshold vs. the number of true features $p_{0}$.
Figure 3: Boxplots comparing Air-HOLP to Ridge-HOLP (with $r=10$) for different values of $R^{2}$ and $p_{0}$. The red line represents the screening threshold $\lceil n/\log(n)\rceil$. Values below this line indicate simulation runs where all true features are screened in.
@@
In conclusion, both Air-HOLP and Ridge-HOLP outperform SIS in correlated settings. However, Air-HOLP demonstrates a consistent advantage over Ridge-HOLP across various settings of $n$, $p$, $\rho$, $p_{0}$, and $R^{2}$, achieving higher sure screening probabilities and lower sure screening thresholds as demonstrated in figures 1-3. This showcases the effectiveness of Air-HOLP’s adaptive penalty selection in enhancing feature screening performance.

### v2

# Air-HOLP: Adaptive Regularized Feature Screening for High Dimensional Correlated Data

@@ Abstract
Handling high-dimensional datasets presents substantial computational challenges, particularly when the number of features far exceeds the number of observations and when features are highly correlated. A modern approach to mitigate these issues is feature screening. In this work, the High-dimensional Ordinary Least-squares Projection (HOLP) feature screening method is advanced by employing adaptive ridge regularization. The impact of the ridge tuning parameter on the Ridge-HOLP method is examined and Adaptive iterative ridge-HOLP (Air-HOLP) is proposed, a data-adaptive advance to Ridge-HOLP where the ridge-regularization tuning parameter is selected iteratively and optimally for better feature screening performance. The proposed method addresses the challenges of tuning parameter selection in high dimensions by offering a computationally efficient and stable alternative to traditional methods like bootstrapping and cross-validation. Air-HOLP is evaluated using simulated data and a prostate cancer genetic dataset. The empirical results demonstrate that Air-HOLP has improved performance over a large range of simulation settings. We provide R codes implementing the Air-HOLP feature screening method and integrating it into existing feature screening methods that utilize the HOLP formula.
Keywords: Correlation analysis, Dimensionality reduction, Regularization, Ridge regression, Sure screening
@@ 1 Introduction
Feature screening is a dimensionality reduction technique that aims to efficiently eliminate numerous non-true features while retaining all true features. Feature screening is often used prior to model selection to handle the curse of dimensionality (Liu and Li, 2020; Liu et al., 2015; Fan et al., 2009). In this paper, we build on the existing High-dimensional Ordinary Least-squares Projection (HOLP) method as introduced in Wang and Leng (2016), by incorporating adaptive ridge regularization. We analyze the impact of the ridge tuning parameter on Ridge-HOLP (Wang and Leng, 2016) and propose the Adaptive iterative ridge-HOLP (Air-HOLP), a method that utilizes a data-adaptive approach for selecting the ridge tuning parameter, which shows remarkable gains in feature screening performance. Additionally, we provide R codes implementing the Air-HOLP feature screening method and integrating it into the group HOLP (Qiu and Ahn, 2020) method which utilizes the HOLP formula. The codes are made available on GitHub at https://github.com/Logic314/Air-HOLP.git.
@@ 2.2 HOLP and Ridge-HOLP
where $r$ is the ridge tuning parameter and $\mathds{I}_{n}$ denotes the identity matrix.
The Ridge-HOLP formula in Equation (3) is equal to the standard ridge formula but is more efficient when $p>n$ as the computational complexity of Ridge-HOLP is $O(n^{3}+n^{2}p)$, while the standard ridge formula is $O(p^{3}+p^{2}n)$. This is because Ridge-HOLP requires an ($n\times p$) by ($p\times n$) multiplication $O(n^{2}p)$ and an ($n\times n$) matrix inverse $O(n^{3})$, while standard ridge requires a ($p\times n$) by ($n\times p$) multiplication $O(p^{2}n)$ and a ($p\times p$) matrix inverse $O(p^{3})$. When $n$ is close to $p$, Wang and Leng (2016) recommend using Ridge-HOLP with a fixed tuning parameter $r=10$.
The penalty in the Ridge-HOLP formula in Equation (3) also helps to handle multicollinearity. When features are highly correlated, then estimated regression coefficients have large variances (Hoerl and Kennard, 1970), which leads to poor ranking accuracy and thus lower screening performance. The penalty in Ridge-HOLP reduces the estimation sensitivity. However, selecting an appropriate tuning parameter $r$ is crucial to balance the bias-variance trade-off and to maximize the screening performance of Ridge-HOLP.
@@
2.3 Challenges of Tuning Parameter Selection in High Dimensions
Several methods exist to choose the ridge tuning parameter. Common approaches include bootstrapping (Delaney and Chatterjee, 1986), cross-validation (Allen, 1971, 1974) and generalized cross-validation (Golub et al., 1979). However, such sub-sampling in the Ridge-HOLP method is computationally expensive due to the repeated computation of the Ridge-HOLP formula. To address this, a more efficient formula called multiridge was proposed by van de Wiel et al. (2021) for applying fast cross-validation in ridge regression. The multiridge formula avoids redundant computations of the ridge formula when applied to multiple values of the tuning parameter $r$, reducing the heavy computations to be repeated only $k$ times in k-fold cross-validation. However, k-fold cross-validation is non-deterministic and can therefore introduce instability in feature screening as different data splits may lead to different screened features. Although repeated cross-validation can mitigate this instability, it comes at the cost of increased computational burden (Martinez et al., 2011).
To avoid the computational burden and instability associated with sub-sampling, an alternative approach is to use a closed-form formula for selecting the tuning parameter $r$. Examples of this alternative approach include Hura Ahmad et al. (2006); Alkhamisi and Shukur (2007); Batah et al. (2008); Dorugade and Kashid (2010); Hoerl and Kennard (1970); Hoerl et al. (1975); Lawless (1976); Kibria (2003); Nomura (1988). However, the formulas proposed in these papers require $p$ to be smaller than $n$. Cule and De Iorio (2013) addressed this limitation by extending the Hoerl et al. (1975) approach to handle $p>n$. Nevertheless, their method involves computing the matrix $\boldsymbol{X}^{\top}\boldsymbol{X}$ and its eigendecomposition, resulting in a time complexity of $O(p^{3}+p^{2}n)$.
@@ 3 The Air-HOLP method
Air-HOLP is an adaptive iterative variant of Ridge-HOLP which efficiently selects the tuning parameter $r$. We begin by discussing how to ensure the selected $r$ in Ridge-HOLP satisfies the sure screening property. Theorem $3$ of Wang and Leng (2016) and Wang et al. (2021) mentions that for the Ridge-HOLP to achieve sure screening, the tuning parameter $r$ must satisfy $r=o(n^{1-(5/2)\tau-\kappa})$, where $01-7.5\tau-2\kappa-\nu>$ 0, and $\tau,\kappa,\nu>0$. It is worth noting that any $r=cn^{a}$ where $0\leq a\leq 0.5$ and $c\geq 0$ satisfies the required condition on $r$. Thus, to ensure that we achieve the sure screening property, we confine the tuning parameter to the range $[0,c\sqrt{n}]$ for some positive constant $c$ and choose $c=1000$ if not otherwise mentioned.
We now discuss how to select the tuning parameter. To select an appropriate tuning parameter that balances the trade-off between bias and variance we solve
@@
The choice of $\boldsymbol{y}_{0}$ in Equation (4) is to circumvent the overfitting that occurs when using $\boldsymbol{y}$ instead. If $\boldsymbol{y}$ was used instead of $\boldsymbol{y}_{0}$, the solution would always be $\hat{r}=0$ because $\boldsymbol{\hat{y}}_{r}=\boldsymbol{y}$ for $r=0$ and $p\geq n$. We evade fitting the noise in $\boldsymbol{y}$ by using $\boldsymbol{y}_{0}$ which represents the noise-free signal. The challenge in solving Equation (5) lies in deriving a computationally efficient and accurate estimate of the unknown expected response $\boldsymbol{y}_{0}$. We address this by proposing the following approach:
  • First, choose an initial tuning parameter $r_{0}$ and use it to compute the initial Ridge-HOLP estimator $\boldsymbol{\hat{\beta}}_{r_{0}}=\boldsymbol{X}^{\top}(\boldsymbol{X}\boldsymbol{X}^{\top}+r_{0}\mathds{I}_{n})^{-1}\boldsymbol{y}$.
  • Then, rank the features in $\boldsymbol{X}$ by the absolute values of their estimated coefficients $|\boldsymbol{\hat{\beta}}_{r_{0}}|$ and screen the top $m^{\prime}$ features, where $m^{\prime}<n$.
  • Finally, fit a regression model utilizing the screened features, and compute the fitted response $\boldsymbol{\tilde{y}}_{r_{0}}$, which serves as the estimate for $\boldsymbol{y}_{0}$.
Notably, the proposed process does not detail a specific method for fitting the model using the screened features. For simplicity, we employ ordinary least squares regression in our empirical research below. Moreover, we use the initial tuning parameter $r_{0}=10$ following Wang and Leng (2016) and choose $m^{\prime}=\lceil n/\log(n)\rceil$ unless otherwise mentioned. Note that $m^{\prime}$ needs to be smaller than $n$ to facilitating fitting an ordinary least squares regression model.
Once $\boldsymbol{y}_{0}$ is estimated, the initial tuning parameter is updated, and the process is repeated iteratively. A given tuning parameter $r_{i}$ is updated through
@@
Air-HOLP iteratively updates the tuning parameter $r$ until $|r_{i+1}-r_{i}|<\delta r_{i+1}$ for some small $\delta>0$, i.e. until the absolute relative error is less than $\delta$ or until a predetermined maximum number of iterations is reached.
@@
$\begin{split}r_{i+1}=\operatorname*{arg\,min}_{0\leq r\leq c\sqrt{n}}{\boldsymbol{y}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}^{2}(\boldsymbol{D}_{n}+r\mathds{I}_{n})^{-2}\boldsymbol{U}^{\top}\boldsymbol{y}-2\boldsymbol{\tilde{y}}_{r_{i}}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}(\boldsymbol{D}_{n}+r\mathds{I}_{n})^{-1}\boldsymbol{U}^{\top}\boldsymbol{y}}.\end{split}$ (8)
Equation (8) can be solved by equating the derivative to 00 as follows:
$\begin{split}-2\boldsymbol{y}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}^{2}(\boldsymbol{D}_{n}+r_{i+1}\mathds{I}_{n})^{-3}\boldsymbol{U}^{\top}\boldsymbol{y}+2\boldsymbol{\tilde{y}}_{r_{i}}^{\top}\boldsymbol{U}\boldsymbol{D}_{n}(\boldsymbol{D}_{n}+r_{i+1}\mathds{I}_{n})^{-2}\boldsymbol{U}^{\top}\boldsymbol{y}=0.\end{split}$ (9)
The roots of Equation (9) may not have a simple closed form solution. Therefore, we use Newton’s method to solve Equation (9).
Algorithm 1 below outlines and summarizes the full process of selecting the tuning parameter $r$ in the Air-HOLP method. Unless otherwise mentioned, the default inputs for Algorithm 1 that we use in our empirical research are $r_{0}=10,m^{\prime}=\lceil n/\log(n)\rceil,c=1000,\delta=0.01,$ and maximum number of iterations $q_{max}=10$.
Algorithm 1 Selection of the ridge tuning parameter $r$
$\boldsymbol{X}$, $\boldsymbol{y}$, $r_{0}$, $m^{\prime}$, c, $\delta,q_{max}$
$\hat{r}$
Step 1: Initialize $i=0\rightarrow r_{i}=r_{0}$.
Step 2: Compute the Ridge-HOLP estimator $\boldsymbol{\hat{\beta}}_{r_{i}}=\boldsymbol{X}^{\top}(\boldsymbol{X}\boldsymbol{X}^{\top}+r_{i}\mathds{I}_{n})^{-1}\boldsymbol{y}$.
Step 3: Rank the features in $\boldsymbol{X}$ by the absolute values of their estimated coefficients $|\boldsymbol{\hat{\beta}}_{r_{i}}|$. Then screen the top $m^{\prime}$ features.
Step 4: Fit an ordinary least squares regression model utilizing the screened features, and compute the fitted response $\boldsymbol{\tilde{y}}_{r_{i}}$.
Step 5: Solve Equation (9) with Newton’s method. If the solution is greater than $c\sqrt{n}$, set $r_{i+1}=c\sqrt{n}$.
Step 6: If $|r_{i+1}-r_{i}|<\delta r_{i+1}$ or if $i\geq q_{max}$, then output $r_{i+1}$. Otherwise, set $i=i+1$ and repeat Steps 2 to 6.
Once the tuning parameter is selected by Algorithm 1, the Air-HOLP method computes the final Ridge-HOLP estimator $\boldsymbol{\hat{\beta}}_{\hat{r}}=\boldsymbol{X}^{\top}(\boldsymbol{X}\boldsymbol{X}^{\top}+\hat{r}\mathds{I}_{n})^{-1}\boldsymbol{y}$, and ranks the features in $\boldsymbol{X}$ by the absolute values of their estimated coefficients $|\boldsymbol{\hat{\beta}}_{\hat{r}}|$. Then screen the top $m$ features. Note that $m$ does not need to be the same as $m^{\prime}$ in Algorithm 1. Both $m$ and $m^{\prime}$ are screening thresholds but they serve different purposes. The threshold $m^{\prime}$ is used to efficiently estimate $\mathbf{y}_{0}$ which leads to a suitable and adaptive selection of the tuning parameter $r$. In contrast, the threshold $m$ is applied later with the goal of eliminating irrelevant features while retaining all true features. One distinction is that $m^{\prime}$ is required to be less than $n$ while $m$ is not.
@@ 4 Simulation Study
In this section, we empirically demonstrate the good performance of Air-HOLP. We primarily compare Air-HOLP’s screening performance to Ridge-HOLP with fixed tuning parameter $r=10$ as recommended by Wang and Leng (2016). While we initially also include Sure Independence Screening (SIS) in our comparison, our main focus is to showcase how Air-HOLP advances Ridge-HOLP by using an adaptive tuning parameter in the more general case when features are correlated. In all simulations, Air-HOLP is implemented with an initial tuning parameter $r_{0}=10$, selected tuning parameter $\hat{r}\in[0,1000\sqrt{n}]$, and screening threshold $m=\lceil n/\log(n)\rceil$. The full simulation results and the code implementing Air-HOLP are both made available on GitHub at https://github.com/Logic314/Air-HOLP.git.
@@ 4.1 Simulation Setup
We generate $5,600$ unique $\boldsymbol{X}$ samples across $112$ distinct simulation settings, with $50$ samples per setting. For each distinct $\boldsymbol{X}$, we generate $250$ $\boldsymbol{y}$’s across $25$ distinct simulation settings. This generates $1,400,000$ datasets [$\boldsymbol{y},\boldsymbol{X}$], encompassing $2,800$ distinct simulation settings, with $500$ samples per setting. The specific simulation settings employed are as follows.
We generate the design matrix $\boldsymbol{X}$ of size $n\times p$ from a multivariate normal distribution with covariance matrix $\boldsymbol{\Sigma}=(1-\rho)\mathds{I}_{p}+\rho$ for varying parameters $n$, $p$, and $\rho$.
@@
In addition to the compound symmetry correlation structure above, we also utilize a spatial correlation structure where only the middle $20\%$ of the features are correlated while the remaining features are independent. Specifically, the covariance matrix of the middle $20\%$ of the features is given by $\boldsymbol{\Sigma}=(1-\rho)\mathds{I}_{0.2p}+\rho$. The parameters $n$, $p$, and $\rho$ vary similarly in both cases of compound symmetry and spatial correlation.
@@ 4.3 Results
Here we analyze the results from all $2,800$ simulation settings to assess the overall performance and comparison of the competing methods. Our analysis begins by demonstrating the effect of correlation on Air-HOLP, Ridge-HOLP, and SIS. Then, we focus on comparing Air-HOLP to Ridge-HOLP aiming to provide a fair and balanced narrative of the overall performance and comparison between them by showing results for representative settings that highlight differences but we conciously refrained from showing settings that give maximal difference between the methods.
Setting 1 - Compound symmetry
Setting 2 - Compound symmetry
Setting 1 - Spatial correlation
Setting 2 - Spatial correlation
Figure 1: Line plots comparing the performances of Air-HOLP, Ridge-HOLP and SIS for different levels of correlation in four different simulation settings. For Ridge-HOLP, four values for the tuning parameter are compared: $r=10,50,200,600$.
@@
In addition, Figure 1 also demonstrates the advantage of adaptive regularization by showcasing four settings were the optimal tuning parameter differ significantly. The smaller tuning parameters ($r=10$ and $r=50$) perform well in Figures 1(a) and 1(c) but not as well in Figures 1(b) and 1(d), whereas the larger tuning parameters ($r=200$ and $r=600$) show the opposite trend. Air-HOLP, on the other hand, consistently performs well in all four settings. On the other hand, when comparing the compound symmetry settings (Figures 1(a) and 1(b)) to the spatial correlation settings (Figures 1(c) and 1(d)) we find that both Air-HOLP and Ridge-HOLP perform better when all features are highly correlated ($\rho=0.9$) compared to when only the middle $20\%$ are correlated. This suggests that the ridge penalty is more effective when the correlations between the features are at a similar level.
Figure 2: Sure Screening Probability Heatmaps. The left panel shows Air-HOLP, the middle panel shows Ridge-HOLP (with $r=10$), and the right panel shows the difference between Air-HOLP and Ridge-HOLP (Sure Screening Probability of Air-HOLP minus Ridge-HOLP). Results shown are for the compound symmetry correlation settings.
We now shift the focus to comparing Air-HOLP and Ridge-HOLP with the recommended tuning parameter of $r=10$, as suggested by Wang and Leng (2016). We mainly focus on the compound symmetry simulation results. However, the same conclusions and insights we discuss apply to the spatial correlation case. Across all $1600$ compound symmetry simulation settings, the difference in sure screening probability between Air-HOLP and Ridge-HOLP ranged from $-0.02$ to $0.60$. In $47\%$ of the settings, the sure screening probabilities for Air-HOLP and Ridge-HOLP were equal, typically occurring when both probabilities were either 00 or $1$. In $45\%$ of the settings, Air-HOLP’s sure screening probability was higher than Ridge-HOLP’s, while in $8\%$ of the settings, it was lower. To better understand the factors influencing this gap in performance, we analyze the impact of $n,p,p_{0}$, and $R^{2}$ on feature screening performance of Air-HOLP and Ridge-HOLP.
@@
(a) The logarithm of the Sure Screening Threshold vs. Theoretical $R^{2}$
(b) The logarithm of the Sure Screening Threshold vs. the number of true features $p_{0}$
Figure 3: Boxplots comparing Air-HOLP to Ridge-HOLP (with $r=10$) for different values of $R^{2}$ and $p_{0}$. The dashed line represents the screening threshold $\lceil n/\log(n)\rceil$. Values below this line indicate simulation runs where all true features are screened in. Results shown are for the compound symmetry correlation settings.
@@
In conclusion, both Air-HOLP and Ridge-HOLP outperform SIS in correlated settings. However, Air-HOLP demonstrates a consistent advantage over Ridge-HOLP across various settings of $n$, $p$, $\rho$, $p_{0}$, and $R^{2}$, achieving higher sure screening probabilities and lower sure screening thresholds as demonstrated in figures 1-3. This showcases the effectiveness of Air-HOLP’s adaptive tuning parameter selection in enhancing feature screening performance.